# Directional Convergence, Benign Overfitting of Gradient Descent in leaky ReLU two-layer Neural Networks

### v1

@@ Appendix A Proof of Directional Convergence
×[$\quad\times\big{[}\left\{\left(q_{-}+\gamma^{2}q_{+}-\gamma\right)n_{-}\|\boldsymbol{\mu}\|^{2}+(q_{-}+\gamma^{2}q_{+})R\right\}n_{+}$
+$\quad+\left\{\left(q_{+}+\gamma^{2}q_{-}-\gamma\right)n_{+}\|\boldsymbol{\mu}\|^{2}+(q_{+}+\gamma^{2}q_{-})R\right\}n_{-}\big{]}$
$\leq\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}\big{\{}(1-\gamma)^{2}n_{+}n_{-}\|\boldsymbol{\mu}\|^{2}+nR\big{\}}$
$\leq\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}\left\{(1-\gamma)^{2}\frac{n^{2}}{4}\|\boldsymbol{\mu}\|^{2}+nR\right\},$
where the last line follows from (33).
Under the regime (ii), we first note
$d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})$
$=$ $(1-\gamma^{2})^{2}q_{+}q_{-}n_{+}n_{-}\|\boldsymbol{\mu}\|^{4}+(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\{n\|\boldsymbol{\mu}\|^{2}+R\}R$
$\geq$ $(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})R^{2}$
$\geq$ $q_{\gamma}^{2}R^{2}.$
With this and (60), we have
$(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\boldsymbol{y}$ $\leq(q_{\gamma}^{2}R^{2})^{-1}\left\{(1-\gamma)^{2}\frac{n^{2}}{4}\|\boldsymbol{\mu}\|^{2}+nR\right\}$ (61)
$\leq q_{\gamma}^{-2}\left\{1+\frac{(1-\gamma)^{2}}{4}\lambda\right\}\frac{n}{R}.$
Therefore, by (57), we have
@@
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$ $\leq q_{\gamma}^{-2}\left\{1+\frac{(1-\gamma)^{2}}{4}\lambda+2\tilde{\varepsilon}\right\}\frac{n}{R}$ (62)
$\leq q_{\gamma}^{-2}\left\{2+\frac{(1-\gamma)^{2}}{4}\lambda\right\}\frac{n}{R}$
$\leq q_{\gamma}^{-2}\left\{2+\frac{1-\gamma}{4\gamma}q_{\gamma}\right\}\frac{n}{R}$
$\leq\frac{3}{\gamma q_{\gamma}^{2}}\frac{n}{R},$
where the second inequality is due to the condition on $\tilde{\varepsilon}$ of the regime (ii) and the third inequality is due to the upper bound on $\lambda$ of the regime (ii).
Under the regime (iii), we have
$d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\geq(1-\gamma^{2})^{2}q_{+}q_{-}n^{2}\|\boldsymbol{\mu}\|^{4}\geq(1-\gamma^{2})^{2}q_{+}q_{-}(1-\tilde{\varepsilon}_{3}^{2})\frac{n^{2}\|\boldsymbol{\mu}\|^{4}}{4},$
where the last inequality is due to (33).
With this and (60), we have
$(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\boldsymbol{y}$ (63)
$\leq$ $\left\{(1-\gamma^{2})q_{+}q_{-}(1-\tilde{\varepsilon}_{3}^{2})\frac{n^{2}\|\boldsymbol{\mu}\|^{4}}{4}\right\}^{-1}\left\{(1-\gamma)^{2}\frac{n^{2}}{4}\|\boldsymbol{\mu}\|^{2}+nR\right\}$
$\leq$ $\frac{(1-\gamma)^{2}+4\lambda^{-1}}{(1-\gamma^{2})^{2}q_{+}q_{-}(1-\tilde{\varepsilon}_{3}^{2})}\|\boldsymbol{\mu}\|^{-2}.$
Therefore, by (57), $\tilde{\varepsilon}_{3}\leq\tfrac{1}{2}$, and $\lambda\geq\tfrac{\gamma}{4}$, we have
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$ (64)
$\leq$ $\frac{(1-\gamma)^{2}+4\lambda^{-1}}{(1-\gamma^{2})^{2}q_{+}q_{-}(1-\tilde{\varepsilon}_{3}^{2})}\|\boldsymbol{\mu}\|^{-2}+\frac{2q_{\gamma}^{-2}n\tilde{\varepsilon}}{R}$
@@
$\leq$ $\left[\frac{(1-\gamma)^{2}+4\lambda^{-1}}{(1-\gamma^{2})^{2}q_{+}q_{-}(1-\tilde{\varepsilon}_{3}^{2})}+\frac{(q_{\gamma}-\gamma)(1-\tilde{\varepsilon}_{3})}{4\big{\{}\frac{(1-\gamma^{2})^{2}}{16}+q_{\gamma}(1+\lambda^{-1})\lambda^{-1}\big{\}}}\frac{2}{\sqrt{n}}\right]\|\boldsymbol{\mu}\|^{-2}$
$\leq$ $\left[\frac{4(1-\gamma)^{2}+64\gamma^{-1}}{3(1-\gamma^{2})^{2}q_{+}q_{-}}+\frac{4(q_{\gamma}-\gamma)}{(1-\gamma^{2})^{2}}\right]\|\boldsymbol{\mu}\|^{-2},$
where the last inequality is due to the condition on $\tilde{\varepsilon}$ of the regime (iii).
Finally, the desired conclusion holds by noting the following inequality
$\frac{n}{n\|\boldsymbol{\mu}\|^{2}+R}\geq\begin{cases}(1+\lambda)^{-1}\frac{n}{R}&\text{if $n\|\boldsymbol{\mu}\|^{2}\leq\lambda R$},\\ (1+\lambda^{-1})^{-1}\|\boldsymbol{\mu}\|^{-2}&\text{if $n\|\boldsymbol{\mu}\|^{2}\geq\lambda R$}.\end{cases}$
We now turn to a lower bound of $\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$.
Lemma B.9.
  1. (i)
    $q_{+}q_{-}=0$, $n\|\boldsymbol{\mu}\|^{2}<\frac{\gamma}{1+(1+\gamma)\tilde{\varepsilon}_{3}}R$, $\tilde{\varepsilon}\leq\min\{\frac{\gamma^{3}}{4\sqrt{n}},\frac{\gamma^{5}}{16(1+\gamma)}\}$, and $\tilde{\varepsilon}_{2}\leq\frac{3\gamma^{2}}{16(1+\gamma)}\frac{\sqrt{n}\|\boldsymbol{\mu}\|^{2}}{R}$,
  2. (ii)
    $q_{+}q_{-}\neq 0$, $n\|\boldsymbol{\mu}\|^{2}\leq\lambda R$, and $\tilde{\varepsilon}\leq\frac{q_{\gamma}^{2}}{4\big{\{}\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\big{\}}^{2}}\min\{\frac{1}{\sqrt{n}},\frac{q_{\gamma}}{4}\}$, $\tilde{\varepsilon}_{2}\leq\frac{3q_{\gamma}^{2}}{16\big{\{}\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\big{\}}^{2}}\frac{\sqrt{n}\|\boldsymbol{\mu}\|^{2}}{R}$ for some $\lambda\leq\frac{q_{\gamma}}{\gamma(1-\gamma)(1+\tilde{\varepsilon}_{3})}$,
  3. (iii)
@@
    $q_{\gamma}>\gamma$, $n\|\boldsymbol{\mu}\|^{2}\geq\lambda R$, $\tilde{\varepsilon}\leq\min\left\{\frac{1-\tilde{\varepsilon}_{3}}{4\sqrt{n}},\frac{q_{\gamma}}{8}\right\}q_{\gamma}^{2}\tilde{C}(\gamma,q_{+},\lambda)\frac{R}{n\|\boldsymbol{\mu}\|^{2}}$, $\tilde{\varepsilon}_{2}\leq\frac{3q_{\gamma}^{2}}{8\sqrt{n}}\tilde{C}(\gamma,q_{+},\lambda)$ for some $\lambda\geq\tfrac{\gamma}{4}$, and $\tilde{\varepsilon}_{3}\leq\tfrac{1}{2}$.
Then, there exists a constant $\tilde{c}_{2}$ which depends only on $q_{+}$ and $\gamma$ such that we have
$\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle\geq\tilde{c}_{2}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}.$
@@
By (56) and Lemma B.5, we have
$\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$ $\geq(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)(\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}+Z\boldsymbol{\mu})$ (65)
$\qquad-\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}\|B^{-1}\boldsymbol{y}\|\|B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)(\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}+Z\boldsymbol{\mu})\|$
$\geq\|\boldsymbol{\mu}\|^{2}(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)\boldsymbol{y}$
$\qquad-R^{-1}\|B^{-1}\boldsymbol{y}\|\|B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)Z\boldsymbol{\mu}\|$
$\qquad\qquad-\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}\|B^{-1}\boldsymbol{y}\|\|B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)(\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}+Z\boldsymbol{\mu})\|$
$\geq\|\boldsymbol{\mu}\|^{2}(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)\boldsymbol{y}$

### v2

@@ Appendix B Proof of Directional Convergence
$\quad\times\Big[(q_{+}+\gamma q_{-})\big\{(q_{-}+\gamma^{2}q_{+}-\gamma)n_{-}\|\boldsymbol{\mu}\|^{2}+(q_{-}+\gamma^{2}q_{+})R\big\}n_{+}$
$\qquad+(q_{-}+\gamma q_{+})\big\{(q_{+}+\gamma^{2}q_{-}-\gamma)n_{+}\|\boldsymbol{\mu}\|^{2}+(q_{+}+\gamma^{2}q_{-})R\big\}n_{-}\Big].$
Thus, we have with Lemma C.7
$(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}(q_{+}B_{+}+q_{-}B_{-})\boldsymbol{y}$ $\leq\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}n(n\|\boldsymbol{\mu}\|^{2}+R)$ (76)
$\leq\frac{8}{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}\frac{n}{n\|\boldsymbol{\mu}\|^{2}+R}.$
Similarly, (35) implies
[…]
$=$ $\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}$
$\qquad\times\Big[(q_{+}+\gamma q_{-})^{2}\big\{(q_{-}+\gamma^{2}q_{+}-\gamma)n_{-}\|\boldsymbol{\mu}\|^{2}+(q_{-}+\gamma^{2}q_{+})R\big\}^{2}n_{+}$
$\qquad\qquad+(q_{-}+\gamma q_{+})^{2}\big\{(q_{+}+\gamma^{2}q_{-}-\gamma)n_{+}\|\boldsymbol{\mu}\|^{2}+(q_{+}+\gamma^{2}q_{-})R\big\}^{2}n_{-}\Big]^{\frac{1}{2}}$
Thus, we have with Lemma C.7
[…]
$\leq\frac{8}{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}\frac{\sqrt{n}}{n\|\boldsymbol{\mu}\|^{2}+R}.$
We now consider regime (ii) and (iii) separately.
Regime (ii): By (66) and (76), we have
@@
$K_{1}\leq\left\{\frac{8}{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}+2(1+\lambda)q_{\gamma}^{-2}\tilde{\varepsilon}\right\}^{2}\frac{n^{2}}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}.$ (79)
Similarly, (67) and (78) imply
@@
$K_{2}\leq\left\{\frac{8}{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}+2(1+\lambda)q_{\gamma}^{-2}\tilde{\varepsilon}\right\}^{2}\frac{n}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}.$ (80)
(54) and (75) imply
$(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}(q_{+}B_{+}+q_{-}B_{-})\boldsymbol{y}$ (81)
$\geq$ $\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}\frac{q_{\gamma}}{2}Rn$
$\geq$ $\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}\frac{q_{\gamma}^{2}}{2(1+\lambda)}n(n\|\boldsymbol{\mu}\|^{2}+R).$
By Lemma C.7, (66), (81), and the assumption on $\tilde{\varepsilon}$, we have for sufficiently large $n$
@@
$K_{1}\geq\left\{\frac{2q_{\gamma}}{5(1+\lambda)}-2(1+\lambda)q_{\gamma}^{-2}\tilde{\varepsilon}\right\}^{2}\frac{n^{2}}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}\geq\left\{\frac{q_{\gamma}}{5(1+\lambda)}\right\}^{2}\frac{n^{2}}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}.$ (82)
Similarly, (54) and (77) imply
[…]
$\geq$ $\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}\left[q_{\gamma}^{2}\left\{\frac{q_{\gamma}}{2}R\right\}^{2}n\right]^{\frac{1}{2}}$
$=$ $\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}\frac{q_{\gamma}^{2}}{2}\sqrt{n}R$
$\geq$ $\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}\frac{q_{\gamma}^{2}}{2(1+\lambda)}\sqrt{n}(n\|\boldsymbol{\mu}\|^{2}+R).$
By Lemma C.7, (67), (83), and the assumption on $\tilde{\varepsilon}$, we have for sufficiently large $n$
$K_{2}\geq\left\{\frac{2q_{\gamma}}{5(1+\lambda)}-2(1+\lambda)q_{\gamma}^{-2}\tilde{\varepsilon}\right\}^{2}\frac{n}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}\geq\left\{\frac{q_{\gamma}}{5(1+\lambda)}\right\}^{2}\frac{n^{2}}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}.$ (84)
By Lemma C.9, (79), (80), (82), and (84), we have
$\left[\left\{\frac{q_{\gamma}}{5(1+\lambda)}\right\}^{2}-\left\{\frac{8}{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}+2q_{\gamma}^{-2}\tilde{\varepsilon}\right\}^{2}\tilde{\varepsilon}\right]\frac{n}{n\|\boldsymbol{\mu}\|^{2}+R}$ (85)
$\quad\qquad\qquad\qquad\leq\|\boldsymbol{\bar{w}}\|^{2}\leq(1+\tilde{\varepsilon})\left\{\frac{8}{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}+2q_{\gamma}^{-2}\tilde{\varepsilon}\right\}^{2}\frac{n}{n\|\boldsymbol{\mu}\|^{2}+R},$
where the left hand side is positive for sufficiently large $n$ due to the assumption on $\tilde{\varepsilon}$.
Regime (iii): By $n\|\boldsymbol{\mu}\|^{2}\geq\lambda R$, $\sqrt{n}\tilde{\varepsilon}\leq\tilde{C}\frac{R}{n\|\boldsymbol{\mu}\|^{2}}\leq\frac{\tilde{C}}{\lambda}$, (66), and (76), we have
$K_{1}$ $\leq\left\{\frac{8}{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}+2q_{\gamma}^{-2}\left(1+\frac{n\|\boldsymbol{\mu}\|^{2}}{R}\right)\tilde{\varepsilon}\right\}^{2}\frac{n^{2}}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}$ (86)
$\leq\left\{\frac{8}{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}+2q_{\gamma}^{-2}\tilde{C}\frac{\lambda^{-1}+1}{\sqrt{n}}\right\}^{2}\frac{n^{2}}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}.$
Similarly, $n\|\boldsymbol{\mu}\|^{2}\geq\lambda R$, $\sqrt{n}\tilde{\varepsilon}\leq\tilde{C}\frac{R}{n\|\boldsymbol{\mu}\|^{2}}\leq\frac{\tilde{C}}{\lambda}$, (67), and (78) imply
$K_{2}$ $\leq\left\{\frac{8}{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}+\left(1+\frac{n\|\boldsymbol{\mu}\|^{2}}{R}\right)\tilde{\varepsilon}\right\}^{2}\frac{n}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}$ (87)
$\leq\left\{\frac{8}{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}+2q_{\gamma}^{-2}\tilde{C}\frac{\lambda^{-1}+1}{\sqrt{n}}\right\}^{2}\frac{n}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}.$
By (42) and (75), we have
$(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}(q_{+}B_{+}+q_{-}B_{-})\boldsymbol{y}$ (88)
$\geq$ $\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}q_{\gamma}(q_{\gamma}-\gamma)\frac{1-\tilde{\varepsilon}_{3}}{2}n(n_{+}+n_{-})\|\boldsymbol{\mu}\|^{2}$
$\geq$ $\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}\frac{(1-\tilde{\varepsilon}_{3})q_{\gamma}(q_{\gamma}-\gamma)}{2}n^{2}\|\boldsymbol{\mu}\|^{2}.$
By Lemma C.7, (66), and (88), we have for sufficiently large $n$
$K_{1}$ $\geq\left\{\frac{2(1-\tilde{\varepsilon}_{3})(q_{\gamma}-\gamma)}{5}\frac{n^{2}\|\boldsymbol{\mu}\|^{2}}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}-2q_{\gamma}^{-2}\left(1+\frac{n\|\boldsymbol{\mu}\|^{2}}{R}\right)\tilde{\varepsilon}\frac{n}{n\|\boldsymbol{\mu}\|^{2}+R}\right\}^{2}$ (89)
$\geq\left\{\frac{2(1-\tilde{\varepsilon}_{3})(q_{\gamma}-\gamma)}{5(1+\lambda^{-1})}-2q_{\gamma}^{-2}\tilde{C}\frac{\lambda^{-1}+1}{\sqrt{n}}\right\}^{2}\frac{n^{2}}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}$
$\geq\left\{\frac{(1-\tilde{\varepsilon}_{3})(q_{\gamma}-\gamma)}{5(1+\lambda^{-1})}\right\}^{2}\frac{n^{2}}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}.$
Similarly, (42) and (77) imply
[…]
$\geq$ $\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}\big[q_{\gamma}^{2}(q_{\gamma}-\gamma)^{2}n_{+}n_{-}(n_{+}+n_{-})\|\boldsymbol{\mu}\|^{4}\big]^{\frac{1}{2}}$
$\geq$ $\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}q_{\gamma}(q_{\gamma}-\gamma)\sqrt{\frac{(1-\tilde{\varepsilon}_{3}^{2})}{4}n^{3}}\|\boldsymbol{\mu}\|^{2}$
$\geq$ $\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}q_{\gamma}(q_{\gamma}-\gamma)\frac{(1-\tilde{\varepsilon}_{3})}{2}n^{\frac{3}{2}}\|\boldsymbol{\mu}\|^{2}$
By Lemma C.7, (66), and (90), we have for sufficiently large $n$
$K_{2}$ $\geq\left\{\frac{2(1-\tilde{\varepsilon}_{3})(q_{\gamma}-\gamma)}{5}\frac{n^{\frac{3}{2}}\|\boldsymbol{\mu}\|^{2}}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}-2q_{\gamma}^{-2}\left(1+\frac{n\|\boldsymbol{\mu}\|^{2}}{R}\right)\tilde{\varepsilon}\frac{\sqrt{n}}{n\|\boldsymbol{\mu}\|^{2}+R}\right\}^{2}$ (91)
$\geq\left\{\frac{2(1-\tilde{\varepsilon}_{3})(q_{\gamma}-\gamma)}{5(1+\lambda^{-1})}\frac{\sqrt{n}}{n\|\boldsymbol{\mu}\|^{2}+R)}-2q_{\gamma}^{-2}\tilde{C}\frac{\lambda^{-1}+1}{\sqrt{n}}\frac{\sqrt{n}}{n\|\boldsymbol{\mu}\|^{2}+R}\right\}^{2}$
$=\left\{\frac{2(1-\tilde{\varepsilon}_{3})(q_{\gamma}-\gamma)}{5(1+\lambda^{-1})}-2q_{\gamma}^{-2}\tilde{C}\frac{\lambda^{-1}+1}{\sqrt{n}}\right\}^{2}\frac{n}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}$
$\geq\left\{\frac{(1-\tilde{\varepsilon}_{3})(q_{\gamma}-\gamma)}{5(1+\lambda^{-1})}\right\}^{2}\frac{n}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}.$
By Lemma C.9, (79), (80), (89), and (91), we have
$\left[\left\{\frac{(1-\tilde{\varepsilon}_{3})(q_{\gamma}-\gamma)}{5(1+\lambda^{-1})}\right\}^{2}-\left\{\frac{8}{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}+2q_{\gamma}^{-2}\tilde{C}\frac{\lambda^{-1}+1}{\sqrt{n}}\right\}^{2}\tilde{\varepsilon}\right]\frac{n}{n\|\boldsymbol{\mu}\|^{2}+R}$ (92)
$\qquad\qquad\leq\|\boldsymbol{\bar{w}}\|^{2}\leq(1+\tilde{\varepsilon})\left\{\frac{8}{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}+2q_{\gamma}^{-2}\tilde{C}\frac{\lambda^{-1}+1}{\sqrt{n}}\right\}^{2}\frac{n}{n\|\boldsymbol{\mu}\|^{2}+R},$
where the left hand side is positive for sufficiently large $n$.
We now turn to obtain bounds of $\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$.
Lemma C.11.
Suppose that all the assumptions of Lemma C.8 hold with $\tilde{\varepsilon}_{3}\leq\frac{1}{2}$ in regime (ii) and (iii). Further suppose $\tilde{\varepsilon}_{2}\leq\tilde{D}\frac{\sqrt{n}\|\boldsymbol{\mu}\|^{2}}{R}$ in (i) and (ii), where $\tilde{D}$ is a constant which depends only on $\gamma$, $q_{+}$ (and $\lambda$ under regime (ii)). Then, there exist constants $\hat{c}_{2},\tilde{c}_{2},N$ which depend only on $\gamma,q_{+}$ (and $\lambda$ under regime (ii) and (iii)) such that $n\geq N$ imply
$\hat{c}_{2}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\leq\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle\leq\tilde{c}_{2}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}.$
@@
By Lemma C.6 and C.7, we have